OpenStudy (anonymous):

Evaluate the definite integral ∫_0^pi/2 3sinxcosx/(1+3sin^2x)^(1/2) dx

OpenStudy (anonymous):

is it from 0 till pi/2 ?

OpenStudy (anonymous):

yes sorry i'm trying to clear it up now

OpenStudy (anonymous):

and the bottom is sin^2(x) ?

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2}{\frac{3sinxcosx}{\sqrt{1+3\sin^2x}}}\]

OpenStudy (anonymous):

yes i hope that clears it up

OpenStudy (anonymous):

I would try with substitution , u=sinx du=cosxdx than it will be 3u/(1+3u^2)^1/2

OpenStudy (anonymous):

u=1+3sin^2x no?

OpenStudy (anonymous):

maybe that would work too

OpenStudy (anonymous):

but I cannot see that through, but with u=sinx, I think I have it

OpenStudy (anonymous):

3u/(1+3u^2)^1/2 until this I guess it is clear now lets use another sub for 1+3u^2=t dt=6udu so it will be 1/(2*(t)^1/2)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

is this clear?

OpenStudy (anonymous):

not thae last part

OpenStudy (anonymous):

the last line

OpenStudy (anonymous):

the antiderivative of u^-1/2

OpenStudy (anonymous):

3u/(1+3u^2)^1/2 we have this and use t=1+3u^2 so dt=6udu so from the nominator 3udu there will be 1/2dt

OpenStudy (anonymous):

I still have not done any integration

OpenStudy (anonymous):

im following you i see what u did in order to get it do be the same as the numerator u divide by 1/2

OpenStudy (anonymous):

yep and integrationg now we get t^1/2+C

OpenStudy (anonymous):

that is (1+3u^2)^1/2 that is (1+3sin^2x)^1/2

OpenStudy (anonymous):

yes it is correct :-) Just checked it on wolfgramalpha

OpenStudy (anonymous):

and now you just have to put pi/2 and 0 into this and subtract

OpenStudy (anonymous):

hmm okay i see haha lol

OpenStudy (anonymous):

sinpi/2=1 sin0=0 2-1=1

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

you are welcome, I like integrals. Lot of work but they are interesting

OpenStudy (anonymous):

yeah they confuse me though but i have to practice...maybe you can help me with a question i posted earlier...please...

OpenStudy (anonymous):

OK, lets see

OpenStudy (anonymous):

which one, you posted quite a few :-)

OpenStudy (anonymous):

lol i think i may have deleted it let me repost